# HoneySat: A Network-based Satellite Honeypot Framework

### v2

@@ VI-C2 Participants
We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to experienced SmallSat mission operators from previously identified missions. In total, we received responses from 10 satellite operators who have between 1 to 10 years of experience operating satellites and have operated between 1 to 5 unique missions. In terms of demographics, 20% (2/10) of the participants were female and 80% (8/10) male. 30% (3/10) belonged to the 18-24 age group, 40% (4/10) to the 25-34 age group, and 30% (3/10) to the 35-44 age group. In regards to geographic location, 40% (4/10) of participants were located in Europe, 20% (2/10) in North America and 40% (4/10) in South America.
Recruiting participants was challenging due to the rarity and specialized nature of the required expertise. Over a span of four months, we reached out to national and international institutions, as well as private corporations involved in operational satellite missions to identify suitable participants.
@@ VI-C3 Methodology and Key Results
Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation of a satellite mission and not a real one. This is because it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes.
Ground Segment Realism. To understand if HoneySat’s ground segment is realistic, participants interacted with HoneySat by performing different satellite operations (discussed in Sec. II-C), to showcase the ground segment components listed in Table IV. One of these tasks is the pass prediction operation which involves calculating when and where a satellite will be within the communication range of a specific ground station. This operation is performed using the ground station control software. After the participants performed this operation on HoneySat, we asked them to rate their perceived level of realism. 90% of the participants strongly agreed and 10% agreed that the pass prediction operation they performed on HoneySat resembles that of a real mission.
@@
After the participants performed the telemetry download operation, we asked them to rate their perceived level of realism of the telemetry output shown during the operation. 90% of the participants agreed or strongly agreed that the telemetry shown during the live telemetry download operation resembles that of a real mission, with the temperature telemetry achieving a 100% strongly agree response rate with some participants mentioning that the temperature values resemble a normal operation and align with the expected temperature values of a satellite. For the EPS telemetry, 80% strongly agreed, 10% agreed, and the last 10% neither agreed nor disagreed when asked if they perceived the EPS telemetry as realistic. In this case, one participant pointed out that in their specific mission they have more batteries, which is expected as satellite missions are highly diverse as we discussed in Sec. II-C. These results indicate that the telemetry generated by the Satellite Simulator is considered highly realistic by the vast majority of participants.
We also asked the participants to assess the ping test operation. This operation is a diagnostic procedure used to verify the communication link between a ground station and the satellite. It involves sending CSP packets between two CSP nodes, namely the ground segment and the spacecraft. After the participants completed this operation’s hands-on task on HoneySat, we asked them to rate their perceived level of realism. 70% of the participants strongly agreed and 30% agreed that the ping test operation resembles that of a real mission with some participants citing that communication was only possible when the satellite was performing a pass.
The results of the ping test operation indicate that HoneySat’s radio simulator (which controls the communication between the ground and space segments) is considered to realistically simulate the communication of a real satellite mission by all participants.
@@
Overall, our survey results provide convincing evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved.
Key finding Q-2 90% of surveyed experienced satellite operators agreed or strongly agreed with the following statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
@@ VI-D Internet Interaction Experiment
This experiment explores HoneySat’s capabilities to entice external actors to interact with it by deploying our honeypot over the Internet.
Experiment Methodology. We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed the necessary TCP ports used by HoneySat’s exposed network protocols discussed in Sec. IV-C, e.g., Telnet’s port 23.
Initially, we deployed our first HoneySat instance to simulate the SUCHAI-2 SmallSat on the premises of the University of Chile, where the mission’s real ground segment is. This on-premises deployment provides the most convincing environment to adversaries. After coordinating with the University of Chile’s local stakeholders for three months, we deployed HoneySat on an onsite server using a University of Chile IP address starting on July 2024.
Although we went to great lengths to make our deployment as realistic as possible, after a few months, our honeypot did not entice any external actors. Therefore, in January 2025, we deployed four additional HoneySat instances on cloud servers. We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [72], and two simulating DLR’s PIXL-1 CubeSat [73]. We selected these two particular CubeSats because they use the CSP ecosystem (Sec. II-E) and are currently in-orbit and operational. We ensured these additional deployments were believable by configuring their satellite personalities and giving them public IP addresses in the regions belonging to the satellite mission origin, namely Germany and the United States. Table V summarizes all five HoneySat deployment details.
TABLE V: HoneySat Internet deployments’ details.
@@
Cloud ACS3 United States 6
Cloud ACS3 United States 6
@@
TABLE VI: Exposed Telnet interactions received.
@@
ACS3 United States Egypt 3 2 hr
PIXL-1 Germany Sweden (Tor Node) 6 5 min
ACS3 United States France (Tor Node) 5 4 min
ACS3 United States USA 8 3 min
Experiment Results. Three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (show in Table VI) via the exposed Telnet protocol discussed in Sec. IV-C. The four interaction sessions yielded 22 satellite flight software-specific commands. These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized, not just haphazardly sending random commands.
@@
It is worth noting that satellites and space systems in general are far less common compared to CPS such as ICS and UAVs making these interactions collected by HoneySat all the more valuable. Additionally, these interactions were produced by human adversaries and not crawlers. We know this because HoneySat’s exposed Telnet was implemented so that a human had to type “activate” before sending commands effectively filtering out Internet Telnet bots.
@@
Key findings Q-2 HoneySat deceived human adversaries who sent 22 satellite flight software-specific commands. The interactions collected by our HoneySat Internet deployments account for six SPACE-SHIELD matrix techniques.
@@ VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
TABLE VII: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
@@
HoneySat CSP
ACS3
NASA
PIXL-1
DLR
SUCHAI-2
U of Chile
CCSDS
PUS
N/A
Experiment Description. We selected the CCSDS ecosystem because it is a standard protocol suite used by other SmallSats [12]. To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [74], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [75].
Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface (Fig. 8). For the mission control software, we used YAMCS’s built-in Mission Control Software [75]. For the radio simulator, we used RACCOON’s communication application. For both the ground configuration and the logging repository, we again used YAMCS built-in features.
Experiment Results. We successfully extended HoneySat to support the CCSDS ecosystem. The implementation was completed by a graduate student with no prior satellite-related in the span of two weeks. However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation. The only extra implementation that was required was the flight software services which we modified using Rust. Other than that, we reused several HoneySat modules such as the satellite personality and the Satellite Simulator.
In summary, these results provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used standard space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
@@ VI-F Case Study: Hardware-in-the-loop Experiment
Experiment Description. In this experiment we integrated HoneySat with an in-orbit, operational SmallSat mission. To achieve this, we collaborated with Quantum Galactics GmbH [76], an aerospace company that develops hardware subsystems for SmallSats. The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand and respond with sanitized telemetry. Due to confidentiality and security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
Experiment Methodology. In order to achieve the HIL integration, we customized both HoneySat and SmallSat X’s mission. For HoneySat, we modified the Satellite Simulator to support SmallSat X’s mission control software (YAMCS). Specifically, we deployed a proxy server that connected the Satellite Simulator to the SmallSat X’s mission production YAMCS instance. Conversely, the SmallSat X’s environment was customized by Quantum Galactics’ team by deploying a script in their YAMCS instance to complete the integration. Once the integration was completed, we coordinated with Quantum Galactics to send a telecommand from HoneySat during one of SmallSat X’s passes.
Experiment Results. The telecommand successfully reached SmallSat X during one of the passes and HoneySat received the appropriate telemetry. Due to security restrictions, we are not able to disclose further details on the telecommand and the telemetry received. However, the entire communication was initiated by HoneySat which in turn received sanitized telemetry from SmallSat X, effectively closing the loop in the HIL experiment. Additionally, Quantum Galactics provided us with a snippet of decoded radio signal data shown in Fig. 5, which confirms that the experiment was successful. In summary, these results highlight HoneySat’s robust extensibility features and provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved.
@@ VII Discussion and Future Work
HoneySat’s Generalizability. HoneySat was designed to be extensible and modular as we discussed in Sec. IV-A which makes our framework generalizable to additional SmallSat missions. We showed this by extending HoneySat to support the two most widely used SmallSat protocol ecosystems, namely CSP and CCSDS (Sec. VI-E), two flight software, SUCHAI and RACOON (Sec. VI-E), and we customized HoneySat to simulate three different SmallSats (Sec. VI-D).
Additionally, we provide a plethora of customization tools including the Satellite Simulator (Sec IV-D), the satellite personality configuration, the ground configuration, and multiple exposed network protocols (Sec V-B2).
Finally, in Table IX, we provide examples of different extensibility use cases, including high-level steps and the HoneySat modules involved so that researchers can use and extend HoneySat for their own experiments.

### v3

@@ VI-C2 Participants
We conducted the survey via Qualtrics and Zoom videoconferencing and distributed the survey directly to operators from previously identified missions. In total, we received responses from 10 satellite operators who have between 1 to 10 years of experience operating satellites and have operated between 1 to 5 unique missions. In terms of demographics, 20% (2/10) of the participants were female and 80% (8/10) male. 30% (3/10) belonged to the 18-24 age group, 40% (4/10) to the 25-34 age group, and 30% (3/10) to the 35-44 age group. In regards to geographic location, 40% (4/10) of participants were located in Europe, 20% (2/10) in North America and 40% (4/10) in South America.
Recruiting participants was challenging due to the rarity and specialized nature of the required expertise. Over a span of four months, we reached out to national and international institutions, as well as private corporations involved in operational satellite missions to identify suitable participants. Despite these challenges, our sample is diverse, including operators from Europe, North America, and South America, spanning industry, government, and academia, across 5 missions.
@@ VI-C3 Methodology and Key Results
Before describing the results, it is important to emphasize that the participants were informed that the system they interacted with was a simulation. We informed participants for two reasons. First, it was not feasible to synchronize participants’ availability to take the survey with the timing of the real satellite’s pass, which may happen only a few times per day and lasts only a few minutes. Second, claiming to provide access to a real satellite would itself be seen as unrealistic.
Ground Segment Realism. To understand if HoneySat’s ground segment is realistic, participants interacted with HoneySat by performing different satellite operations (discussed in Sec. II-C), to showcase the ground segment components listed in Table III. One of these tasks is the pass prediction operation which involves calculating when and where a satellite will be within the communication range of a specific ground station. This operation is performed using the ground station control software. After the participants performed this operation on HoneySat, we asked them to rate their perceived level of realism. 90% of the participants strongly agreed and 10% agreed that the pass prediction operation they performed on HoneySat resembles that of a real mission.
@@
After the participants performed the telemetry download operation, we asked them to rate their perceived level of realism of the telemetry output shown during the operation. 90% of the participants agreed or strongly agreed that the telemetry shown during the live telemetry download operation resembles that of a real mission, with the temperature telemetry achieving a 100% strongly agree response rate with some participants mentioning that the temperature values resemble a normal operation and align with the expected temperature values of a satellite. These results indicate that the telemetry generated by the Satellite Simulator is considered highly realistic by the vast majority of participants.
@@
Overall, our survey results provide evidence for answering question Q-2 in the affirmative and design objective DO-2 as achieved. These findings serve as empirical evidence of HoneySat’s fidelity when simulating a real satellite mission. Our survey’s questions and results are available online [^3] https://github.com/HoneySat/honeysat-survey-data..
Key finding Q-2 90% of surveyed satellite operators agreed with the statement: “I would not be able to distinguish between HoneySat’s satellite honeypot system from the real CubeSat satellite mission it is based on.”
@@ VI-D Internet Interaction Experiment
This experiment explores HoneySat’s capabilities to entice external actors by deploying it over the Internet.
Experiment Methodology. We leveraged HoneySat’s customization and deployability features to deploy five instances of HoneySat over the Internet and exposed TCP ports 23 for Telnet, 80 for the web interface, and 5901 for VNC.
Deployment. Initially, we deployed our first HoneySat instance to simulate the SUCHAI-2 SmallSat on the premises of the University of Chile, where the mission’s real ground segment is. This on-premises deployment provides the most convincing environment to adversaries. After coordinating with the University of Chile’s local stakeholders for three months, we deployed HoneySat on an onsite server using a University of Chile IP address starting on July 2024.
Although we went to great lengths to make our deployment as realistic as possible, after a few months, our honeypot did not entice any external actors. Therefore, in January 2025, we deployed four additional HoneySat instances on cloud servers. We configured two instances to simulate NASA’s Advanced Composite Solar Sail System (ACS3) CubeSat [nasa2025acs3], and two simulating DLR’s PIXL-1 CubeSat [eoportal2025pixl1]. We selected these two particular CubeSats because they use the CSP ecosystem (Sec. II-E) and are currently in-orbit and operational. We ensured these additional deployments were believable by configuring their satellite personalities and giving them public IP addresses in the regions belonging to the satellite mission origin, namely Germany and the United States. Table IV summarizes all five HoneySat deployment details.
While IP assignment can affect our honeypots’ covert status [lopez2022honeyplc], the increasing use of cloud and web-based ground segment services, such as AWS Ground Station [amazon2024satellite] and YAMCS [yamcs], has made cloud-hosted satellite operations increasingly commonplace. As a result, a cloud deployment no longer inherently signals inauthenticity to adversaries. Furthermore, the part of the Ground Segment that we expose on cloud IPs is not dependent on being collocated with the GS and thus may as well be hosted with a cloud provider.
TABLE IV: HoneySat Internet deployments’ details.
@@
Cloud ACS3 USA 6
Cloud ACS3 USA 6
@@
TABLE V: Exposed Telnet interactions received.
Date
@@
Jan 18, 2025 ACS3 USA Egypt 4 2 hr
Jan 24, 2025 PIXL-1 Germany Sweden (Tor) 6 4 min
Jan 24, 2025 ACS3 USA France (Tor) 9 5 min
Apr 3, 2025 ACS3 USA USA 8 3 min
Experiment Results. We collected several gigabytes of network traffic data on ports 23, 80 and 5901 for each honeypot, however most of these data was generated by crawlers and bots and not satellite-specific. But, on January 18th, 2025, we received our first space-specific interaction via Telnet and in the next few months we received more. Overall, three of our honeypots successfully enticed external actors and captured four distinct interaction sessions (shown in Table V) all via Telnet. The four interaction sessions yielded 22 flight software-specific commands. These flight software-specific commands show that the adversaries that interacted with HoneySat were purposefully trying to exploit our honeypot using commands that the flight software recognized.
Across the four Internet interaction sessions we observed, the first lasted two hours and originated from a non-Tor IP in Egypt. The next two sessions occurred within a week, shared their command set (e.g., help, ls, and fp_show), and transitioned to Tor exit nodes in Sweden and France. These three factors suggest that the same adversary returned using anonymized infrastructure after initial reconnaissance. The first session may have ended after the actor gathered sufficient information. In contrast, the fourth and final interaction in April, separated by over two months, from a non-Tor U.S. address, and using a very different command set appears unrelated and likely originates from a different actor. Overall, the first three interactions are best interpreted as a single adversary adapting its anonymity strategy, whereas the last session was likely initiated by a second actor.
@@
While our dataset is relatively small, its significance becomes evident when contextualized. For example, according to a 2018 Open Platform Communications (OPC) group report [arc2018opcinstalledbase], there are an estimated 47 million OPC-enabled ICS devices deployed worldwide, compared to 11,000 operational satellites in 2025 [livescience2024satellites]. This represents about 4,700 ICS devices for every one satellite. In other words, attackers have orders of magnitude more opportunities to target ICS devices than satellites. Thus, making our dataset a rare and valuable contribution to understanding space systems’ TTPs.
In addition to its rarity, our dataset is the first empirical real-world dataset of TTPs against satellites. As such, it not only provides the first view into how adversaries interact with satellite systems, but also establishes a reproducible baseline reference for future studies. Lastly, our contribution complements prior work that provides limited insights, e.g., no specific telecommands [uscc2011annualreport].
@@
Key findings Q-2 HoneySat deceived human adversaries who sent 22 satellite flight software-specific commands. The interactions collected by our HoneySat deployments comprise six SPACE-SHIELD techniques.
@@ VI-E Case Study: Extending HoneySat to CCSDS Ecosystem
Experiment Description. We selected CCSDS because it is a standard protocol suite used by other SmallSats [willbold2023space]. To accomplish this, we leveraged an open-source CCSDS ecosystem-based flight software framework, RACCOON OS [raccoonos], and YAMCS, an open-source Mission Control software framework with built-in support for PUS [yamcs].
Regarding the ground segment, we implemented the exposed network protocol using YAMCS’ built-in web interface. For the mission control software, we used YAMCS’s built-in Mission Control Software [yamcs]. For the radio simulator, we used RACCOON’s communication application. For both the ground configuration and the logging repository, we again used YAMCS built-in features.
Experiment Results. We successfully extended HoneySat to support the CCSDS ecosystem. The implementation was completed by a graduate student with no prior satellite-related experience in the span of two weeks. However, the majority of the effort involved in extending HoneySat to support the CCSDS ecosystem was dedicated to understanding the ecosystem itself, and analyzing the RACCOON flight software code and the YAMCS framework documentation. The only extra implementation that was required was the flight software services which we modified using Rust. Other than that, we reused several modules such as the Satellite Simulator.
In summary, these results provide evidence for answering Q-3 in the affirmative and DO-3 as achieved.
Key findings Q-3 Out of the box HoneySat supports CSP and CCSDS, the two most widely used space ecosystems, and was evaluated by simulating 3 real-world SmallSats.
TABLE VI: SmallSat ecosystems (2) and example SmallSats (3) integrated to HoneySat.
@@
HoneySat CSP ACS3 NASA
PIXL-1 DLR
SUCHAI-2 U of Chile
CCSDS PUS N/A
@@ VI-F Case Study: Hardware-in-the-loop Experiment
Experiment Description. In this experiment we integrated HoneySat with an in-orbit, operational SmallSat mission. To achieve this, we collaborated with an aerospace company that develops hardware subsystems for SmallSats. The experiment involved sending a telecommand from HoneySat’s ground segment simulation which was then sent to the in-orbit SmallSat, which would then process the telecommand. Due to security restrictions, we are unable to disclose the name of the in-orbit SmallSat which we refer to as SmallSat X.
Experiment Methodology. In order to achieve the HIL integration, we customized both HoneySat and SmallSat X’s mission. For HoneySat, we modified the Satellite Simulator to support SmallSat X’s mission control software (YAMCS). Specifically, we deployed a proxy server that connected the Satellite Simulator to the SmallSat X’s mission production YAMCS instance. Conversely, the SmallSat X’s environment was customized by aerospace company’s team by deploying a script in their YAMCS instance to complete the integration. Once the integration was completed, we coordinated with the aerospace company to send a telecommand from HoneySat during one of SmallSat X’s passes.
Experiment Results. The telecommand successfully reached SmallSat X during one of the passes and HoneySat received the appropriate telemetry. Due to security restrictions, we are not able to disclose details on the telecommand and the telemetry received. However, the entire communication was initiated by HoneySat which in turn received sanitized telemetry from SmallSat X, effectively closing the loop in the HIL experiment. Additionally, the aerospace company provided us with a snippet of radio signal data shown in Fig. 7, which confirms that the experiment was successful. In summary, these results highlight HoneySat’s robust extensibility features and provide strong evidence for answering question Q-3 in the affirmative and design objective DO-3 as achieved. These findings serve as empirical evidence of HoneySat’s fidelity when interacting with a real satellite mission.
@@ VII Discussion and Future Work
Challenges of Creating the First Satellite Honeypot. During HoneySat’s design and implementation we encountered and overcame two main challenges which stemmed from fundamental differences between satellites and other Cyber-Physical Systems (CPS), e.g., ICS.
First, time and link-constrained communication. Unlike ICS honeypots, which assume continuous network connectivity and stable control loops, e.g., scan cycles, satellite communication occurs only during orbital passes with variable duration and packet loss. To overcome this challenge, we developed the Radio Simulator discussed in Sec. IV-C.
Second, dynamic physics-aware simulation. ICS honeypots simulate sensor readings through simple physics loops (e.g., tank level, valve position). In contrast, a satellite honeypot must dynamically maintain physically coherent telemetry across interdependent subsystems such as attitude, and power. We addressed this by creating the Satellite Simulator, which simulates subsystems and sensors that communicate between each other and generate data in real time.
Satellite Honeynets. To increase attacker engagement and reflect emerging satellite networks [ma2023network], multiple HoneySat instances can be deployed as a honeynet [icsnet2024salazar]. Each instance could simulate different satellites and communicate through virtual inter-satellite links (ISLs). In addition, HoneySat could also be part of a honeynet connected to other ground mission infrastructure such as database servers and radio equipment. Due to HoneySat’s extensibility, these honeynet scenarios could be implemented in the future.
Satellite Honeypots’ Fingerprinting. Based on our experience designing, implementing and evaluating HoneySat we now discuss some qualitative detection vectors that can guide future satellite honeypot fingerprinting research.
Space Protocol Uniformity. The use of default space protocol configurations may be used for fingerprinting. For example, CSP deployments built directly from libCSP expose identical port assignments and diagnostic services, e.g. ping.
Perfectly periodic pass timing. Real satellites rarely follow identical pass schedules; orbital perturbations, TLE drift, and operational delays introduce small timing variations. A honeypot that enables communication at perfectly fixed intervals or with constant latency can thus be fingerprinted. To avoid this, HoneySat periodically downloads the latest TLE data to keep the orbital simulation from drifting. This, on the other hand, causes a jump in the orbit once the new TLE data is used to base the simulation on.
Telecommand implementation. A given FS may have completely different telecommand names and functions depending on the satellite and thus a mismatch may reveal the honeypot. Additionally, telecommands may have revealing features that print information about the underlying honeypot host, e.g., Ubuntu, which may reveal the simulation.
VII-A Initial Satellite Honeypot Anti-Detection Techniques
Building upon lessons learned from HoneySat’s development and from other CPS honeypots [uitto2017survey, tay2023taxonomy], we identify concrete strategies to enhance HoneySat’s and future satellite honeypots’ resistance to fingerprinting and detection.
Improve Environmental Diversity
Previous research shows that static structural layouts and repeated host fingerprints enable easy detection [tay2023taxonomy]. Future satellite honeypots and honeynets can incorporate diversity in its ground segment topology, and network services. Following the Purdue-style segmentation [298052] used in ICS honeypots, different HoneySat modules (mission control, station control, payload) can be isolated behind realistic network layers such as routers and firewalls.
Implement Adaptive Reconfiguration
Honeypot adaptive reconfiguration occurs when a honeypot recognizes that it is being probed or scanned, e.g., Nmap probes, and reconfigures itself to avoid detection [gabrys2024honeyganpotsdeeplearning, uitto2017survey]. Adapting this technique to satellite honeypots would involve both the ground and space segments. For example, the honeypot could modify exposed services, e.g., CSP over ZeroMQ to appear as a different ground station node. Additionally, different telemetry datasets could be rotated, or switched between low interaction and high interaction versions of the honeypot [kyung2017honeyproxy].
Use Real hardware and software identifiers
Future satellite honeypots could use real hardware identity and fingerprints such as MAC Organizationally Unique Identifiers (OUIs), and telemetry headers such as APIDs (Application Process Identifiers) [SpacePacketProtocol2020]. Obtaining these details is non-trivial as they are often proprietary or absent from public documentation, and may require reverse-engineering, or careful analysis of captured mission traffic.